# DiffPatch: Generating Customizable Adversarial Patches using Diffusion Model

### v1

@@ 3.3 Target Mask Control
Our method can generate adversarial patches of various shapes, with the process being controlled by the target mask $m$. Using target masks serves not only to extract patches of different shapes but also plays a crucial role in the natural appearance of the patch. Since Stable Diffusion [39] generates square images, there is environmental information surrounding our target. After selecting the initial image $\mathcal{I}$, to prevent the background from interfering with the main subject during the optimization process, we need to replace the background with a solid-colored image $s$ before the Image-to-Latent Mapping stage:
$\mathcal{I}^{\prime}=\mathcal{I}\odot m+s\odot(1-m),$ (11)
@@
Overall, the DiffPatch generation process starts with Image-to-Latent Mapping. Incomplete Diffusion Optimization is then applied to reduce memory consumption and enhance performance. Finally, Target Mask Control is used to maintain a natural appearance. The generation process is illustrated in Algorithm 1. In addition, the patch can be iteratively optimized using DiffPatch to achieve an accumulative effect, as shown in Figure 3. Specifically, the generated patch can be used to generate a new patch, enabling stronger attack performance while maintaining natural appearance.
Figure 3: The effect of iterative generation (accumulated style variation from left to right).
@@ 4 Experiments
[…]
Method YLv3 YLv3t YLv4 YLv4t YLv5s YLv7t FRCNN DETR
AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR
${}^{(P_{A})}$dvYL 31.9 |||| 46.3 13.6 |||| 66.4 46.3 |||| 36.2 9.0 |||| 69.0 24.8 |||| 53.4 61.8 |||| 20.7 38.3 |||| 24.0 33.1 |||| 36.7
${}^{(P_{B})}$-SEA 22.0 |||| 39.0 7.9 |||| 59.0 18.4 |||| 48.7 12.2 |||| 56.8 10.2 |||| 49.3 46.3 |||| 26.6 20.0 |||| 27.1 21.8 |||| 36.3
${}^{(P_{C})}$ 50.4 |||| 33.7 15.9 |||| 74.2 61.3 |||| 26.2 13.7 |||| 78.9 41.0 |||| 41.4 69.0 |||| 17.5 55.5 |||| 25.1 43.3 |||| 41.5
${}^{(P_{D})}$2D 50.3 |||| 33.9 18.8 |||| 69.1 58.4 |||| 30.9 15.6 |||| 75.0 41.0 |||| 40.7 68.7 |||| 18.4 52.9 |||| 25.0 44.8 |||| 34.2
${}^{(P_{1})}$3(Ours) 12.9 |||| 70.4 17.0 |||| 64.1 58.2 |||| 28.1 21.3 |||| 60.4 36.0 |||| 52.2 60.2 |||| 24.6 38.6 |||| 29.3 38.6 |||| 38.6
${}^{(P_{2})}$3t(Ours) 40.2 |||| 38.6 3.9 |||| 86.1 62.5 |||| 24.4 16.9 |||| 65.3 40.6 |||| 44.8 59.4 |||| 24.1 44.1 |||| 29.5 40.4 |||| 39.6
${}^{(P_{3})}$4(Ours) 32.3 |||| 45.0 23.8 |||| 57.5 35.1 |||| 49.9 31.1 |||| 46.8 45.1 |||| 43.5 62.4 |||| 21.6 43.7 |||| 27.4 39.0 |||| 39.0
${}^{(P_{4})}$4t(Ours) 40.8 |||| 41.0 11.8 |||| 73.6 61.6 |||| 25.5 9.2 |||| 79.1 38.3 |||| 50.3 57.4 |||| 25.7 45.3 |||| 28.4 39.5 |||| 41.4
${}^{(P_{5})}$5s(Ours) 33.9 |||| 47.4 17.8 |||| 64.3 60.3 |||| 27.2 23.7 |||| 56.5 22.0 |||| 69.9 61.3 |||| 23.1 40.6 |||| 29.9 37.2 |||| 38.3
${}^{(P_{6})}$7t(Ours) 49.6 |||| 35.0 19.9 |||| 64.3 64.2 |||| 23.0 26.0 |||| 56.7 39.6 |||| 44.5 51.7 |||| 33.5 50.2 |||| 26.1 42.2 |||| 38.9
${}^{(P_{7})}$(Ours) 38.8 |||| 42.7 19.9 |||| 62.4 58.1 |||| 29.2 23.0 |||| 60.8 37.2 |||| 50.7 61.7 |||| 24.3 22.6 |||| 38.5 35.4 |||| 43.7
${}^{(P_{8})}$(Ours) 46.6 |||| 35.2 35.2 |||| 49.0 67.1 |||| 22.1 34.3 |||| 46.7 44.6 |||| 43.1 66.8 |||| 20.5 46.2 |||| 26.1 37.1 |||| 42.0
${}^{(P_{j})}$ource 53.5 |||| 29.7 45.9 |||| 38.1 70.0 |||| 18.5 47.8 |||| 36.9 55.8 |||| 35.0 69.4 |||| 18.4 53.5 |||| 23.3 47.3 |||| 33.6
       Gray 72.3 |||| 16.5 61.8 |||| 25.2 76.8 |||| 12.8 56.4 |||| 30.0 74.1 |||| 15.5 77.1 |||| 13.5 64.4 |||| 17.3 50.2 |||| 34.4
${}^{(P_{m})}$andom 70.2 |||| 16.9 64.3 |||| 22.2 78.5 |||| 12.2 60.4 |||| 26.9 75.1 |||| 15.5 79.8 |||| 12.3 67.3 |||| 17.6 51.1 |||| 31.7
${}^{(P_{n})}$andom 72.8 |||| 17.0 68.0 |||| 19.0 82.4 |||| 10.4 69.8 |||| 20.7 83.2 |||| 11.7 79.3 |||| 12.4 72.0 |||| 16.4 57.0 |||| 25.9
$P_{1}$ $P_{2}$ $P_{3}$ $P_{4}$ $P_{5}$ $P_{6}$ $P_{7}$ $P_{8}$ $P_{A}$ $P_{B}$ $P_{C}$ $P_{D}$ $P_{j}$ $P_{m}$ $P_{n}$
@@ 4.1 Experimental Setup
In our experiments, we trained and evaluated the adversarial patches using the INRIA person dataset [11], which consists of 614 training images and 288 test images. To meet the input requirements of the detector, we padded all images with gray pixels to make them square and then resized them to 640 $\times$ 640. Direct resizing would distort the image proportions. Furthermore, we evaluated the transferability of different adversarial patch methods on the MPII Human Pose dataset [1], COCO dataset [31] and 1,000 images from diverse scenarios collected from the Internet.
Attacked Models
Our experiments attack a variety of detectors, including YOLOv3, YOLOv3-tiny [37], YOLOv4, YOLOv4-tiny [4], YOLOv5s [25], YOLOv7-tiny [44], FasterRCNN [38] and DETR [6]. We will abbreviate YOLO as YL and FasterRCNN as FRCNN for ease of presentation, such as YOLOv4-tiny as YLv4t. All models are pre-trained on COCO dataset [31].
@@
We treat person as the target attack class, set the batch size to 32, the maximum epoch $N_{iter}=200$, DDIM sampling steps $T=50$, unconditional embedding optimization iterations $N_{u}=10$, $\eta_{u}=0.01$, $\eta_{p}=0.003$, $\epsilon=0.5$, we use Adam optimizer [26] with the default parameters, the textual prompts $\mathcal{P}$ are automatically generated using BLIP [29], the version of Stable Diffusion [39] is v1.4. Since we are generating irregular patches, we set the patch size to match the pixel area of a square patch with a scale of $\tau=0.2$. To achieve stronger performance in patch attacks, we applied two iterations of optimization to the patch using DiffPatch. We regard the prediction of detector on clean data as the ground truth, and measure the capability of adversarial patch attack using the Average Precision (AP). A lower AP indicates that the accuracy of the detector in recognizing persons is reduced, indirectly indicating stronger patch attack performance. Additionally, since AP only measures the overall decline trend caused by the patches and does not directly reflect the attack performance, we also evaluated the Attack Success Rate (ASR) of various patches. The confidence threshold and IoU threshold for the detector are both set to 0.5. All of our experiments were conducted on a single NVIDIA Tesla A100.
@@ 4.2 Attack Performance Evaluation
We generated a series of adversarial patches targeting different models using DiffPatch and compared them with current patch generation methods on INRIA dataset, including AdvYL [43], T-SEA [22], NPAP [21], and D2D [30]. The detailed results can be seen in Table 1. DiffPatch ($P_{1}-P_{8}$) performs comparably to the best existing non-natural patch generation methods in terms of the AP metric. Notably, DiffPatch achieved the best Attack Success Rate (ASR), with an improvement of 10.1% over AdvYL on YLv4t. Compared to naturalistic adversarial patches ($P_{C}$ and $P_{D}$), our patch trained on YLv4t effectively reduced the AP by 4.5% and 6.4%, and increased the ASR by 0.2% and 4.1%.
@@ 4.3 Cross-dataset Evaluation
Considering the relatively simple scenes in the INRIA dataset, we conducted tests to evaluate the attack performance of the patches in more diverse environments. We trained patches using YLv4t with various methods and tested them on multiple datasets, including MPII and COCO. Due to the large size of these datasets, we randomly sampled a subset of images containing people for testing, selecting 729 images from COCO and 569 images from MPII. Additionally, since the images from these two datasets are relatively outdated with low-resolution images, we also collected 1,000 high-resolution images from the Internet, representing diverse scenes (denoted as “Collected”). The results, as shown in Table 2, demonstrate that our method significantly reduces the AP while substantially increasing the ASR. On average, the ASR shows a 17.5% improvement over the unnatural patch method T-SEA and a 2.6% improvement over the natural patch method NPAP.
@@
ASR%$\uparrow$ INRIA 56.8 78.9 75.0 79.1
MPII 60.3 73.4 73.3 75.9
COCO 54.5 63.9 60.3 67.8
Collected 67.2 82.1 79.4 85.8
Table 2: Attack performance of different attacks across various datasets.
4.4 Ensemble Attack Evaluation
We employed an ensemble approach, training our patch using the averaged gradients from multiple models. Compared to other natural patch generation methods, our single patch achieved strong attack performance across various object detection models. As shown in Table 3, NPAP was trained on YLv3 and YLv4t, while D2D and our DiffPatch were trained on all models. In white-box attacks, our method demonstrated a significant performance improvement over D2D.
@@ 4.4 Ensemble Attack Evaluation
Method YLv3 YLv3t YLv4 YLv4t YLv5s YLv7t FRCNN Avg.
AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR AP |||| ASR
NPAP 45.2 |||| 34.2 16.2 |||| 74.8 65.8 |||| 22.6 19.3 |||| 71.7 57.4 |||| 28.5 70.8 |||| 17.3 57.8 |||| 24.0 32.3 |||| 53.0
D2D 48.9 |||| 35.7 17.5 |||| 66.8 58.4 |||| 28.9 24.6 |||| 59.3 35.6 |||| 43.2 67.0 |||| 19.5 52.7 |||| 24.3 43.5 |||| 39.7
DiffPatch 9.1 |||| 65.9 6.1 |||| 79.5 28.9 |||| 47.4 12.0 |||| 75.5 23.2 |||| 62.4 55.4 |||| 28.5 19.9 |||| 38.9 22.1 |||| 56.9
@@ 4.5 Ablation Studies
We conducted experiments using both the Common Detection Loss and IoU-Detection Loss under the same patch training setting. As shown in Figure 4, compared to the Common Detection Loss, our IoU-Detection Loss demonstrates greater stability during training process and achieves higher attack performance in a earlier time.
@@
The impact of patch shapes was investigated, as indicated by the data for $P_{m}$ and $P_{n}$ in Table 1. We compared the performance of random noise images with square and irregular shapes (with the same size as DiffPatch). From the AP and ASR metrics, it is evident that square patches yield higher attack benefits compared to irregular patches. In other words, designing irregular patches with high attack success rates is more challenging.
@@
Using latent vectors from intermediate timestamps to generate patches effectively mitigates semantic loss. We tested the results of latent vectors at different timestamps, evaluating the attack performance and similarity to textual descriptions for patches generated from diffusion trajectories of varying lengths. See Figure 5 for details. We use the cosine similarity of CLIP [35] embeddings to measure the alignment between the patch and the prompt, which indirectly reflects the visual quality of the patch. It is evident that starting diffusion optimization from $\frac{T}{2}$ (25 Steps) not only achieves a high success rate but also produces images that better align with textual descriptions, thereby maintaining greater semantic consistency with the original images.
@@
The target mask mechanism helps mitigate semantic loss in images to some extent. We compared the training results with and without the use of this mechanism. Without using target mask leads to quality issues in the patches during optimization, as shown in Figure 6.
Figure 6: The impact of target mask on adversarial patch optimization. Without target mask control, adversarial patches exhibit noticeable detail distortions (indicated by red boxes).
Optimization Constraint
We use the infinity norm to constrain each optimization of the patch to ensure its natural appearance. The size of $\epsilon$ significantly impacts patch generation. An appropriate $\epsilon$ can effectively balance the trade-off between attack performance and appearance. As shown in Table 4, we ultimately chose $\epsilon=0.5$ as the constraint parameter for patch generation, as it maintains both appearance and a high attack success rate.
Constraint $\epsilon$ 0.3 0.5 0.7
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
CLIP Sim. 75.6% 73.2% 63.9%
ASR 57.0% 63.0% 67.4%
Table 4: The effect of optimization constraint $\epsilon$ on latent vectors.
@@
As the results in Table 5, iterative optimization can further enhance the adversarial strength of the patch. However, it also makes the patch unnatural.
@@
4.6 Physical-world Adversarial Patch Dataset
Our DiffPatch achieves promising attack results even in physical-world environments. Using DiffPatch, we designed 9 adversarial patches and printed them on T-shirts, as shown in Figure 7. We recruited 20 participants (with ethics approval) to capture images in diverse indoor and outdoor, including laboratories, campus, cafeteria, subway station, and shopping mall. In total, we collected 1,131 images into our AdvPatch-1K dataset, including both individual and group (2 to 10+ persons) photos. The dataset is accompanied with detailed annotations of the person and patch locations (bounding boxes). Several example images are shown in Figure 1, with detailed statistics in Figure 8. The images in AdvPatch-1K can effectively evade the YOLOv5s detector, as shown in the right subfigure of Figure 8. Our dataset provides a valuable resource for advancing defense research against adversarial patch attacks.
Figure 7: Adversarial T-shirts created using DiffPatch.
Figure 8: Left: The proportion of images with different numbers of persons in AdvPatch-1K. Right: The detection performance of YOLOv5s on AdvPatch-1K under different confidence thresholds.

### v3

@@ 3.3 Target Mask Control
Our method can generate adversarial patches of various shapes, controlled by the target mask $m$. Using target masks not only allows for the extraction of patches of different shapes but also plays a crucial role in maintaining the natural appearance of the patch. Since Stable Diffusion [45] generates square images, there is environmental information surrounding our target. After selecting the initial image $\mathcal{I}$, to prevent the background from interfering with the main subject during the optimization process, we replace the background with a solid-colored image $s$ before the Image-to-Latent Mapping stage:
$\mathcal{I}^{\prime}=\mathcal{I}\odot m+s\odot(1-m).$ (10)
@@
Overall, the DiffPatch generation process begins with Image-to-Latent Mapping, followed by Incomplete Diffusion Optimization to reduce memory consumption and enhance performance. Finally, Target Mask Control is applied to preserve the natural appearance of the generated patches. The complete generation process is outlined in Algorithm 1. Furthermore, the patch can be iteratively optimized using DiffPatch to achieve an accumulative effect, as illustrated in Figure 3. Specifically, the generated patch can serve as the input for generating a new patch, enabling stronger attack performance while maintaining a natural appearance. This iterative approach ensures both effectiveness and stealthiness in adversarial patch generation.
Figure 3: The effect of iterative generation (showing accumulated style variation from left to right).
@@ 4 Experiments
Table 1: Comparing DiffPatch with other methods in ASR (%) (higher is better). The best results are underlined. denotes transfer attack methods, while the rest are white-box attacks. trained on YLv2, trained on YLv4t. The generated patches are shown at the bottom.
Method ASR ($\uparrow$)
YLv3 YLv3t YLv4 YLv4t YLv5s YLv7t YLv10s YLv12s FRCNN DETR RT-DETR
${}^{(P_{A})}$dvYL 46.3 66.4 36.2 69.0 53.4 20.7 35.1 30.6 24.0 36.7 30.6
${}^{(P_{B})}$-SEA 39.0 59.0 48.7 56.8 49.3 26.6 61.1 62.1 27.1 36.3 32.1
${}^{(P_{C})}$ 33.7 74.2 26.2 78.9 41.4 17.5 33.4 22.0 25.1 41.5 48.3
${}^{(P_{D})}$2D 33.9 69.1 30.9 75.0 40.7 18.4 23.1 26.0 25.0 34.2 39.7
${}^{(P_{E})}$ 39.0 49.5 32.9 46.0 43 18.3 20.7 23.0 22.9 39.8 25.6
${}^{(P_{F})}$dvART 22.9 36.4 15.4 33.9 19.3 17.3 12.1 11.4 20.4 31.0 19.0
${}^{(P_{1})}$v3(Ours) 70.4 64.1 28.1 60.4 52.2 24.6 26.6 26.8 29.3 38.6 44.9
${}^{(P_{2})}$v3t(Ours) 38.6 86.1 24.4 65.3 44.8 24.1 24.0 19.9 29.5 39.6 46.7
${}^{(P_{3})}$v4(Ours) 45.0 57.5 49.9 46.8 43.5 21.6 27.5 24.8 27.4 39.0 45.2
${}^{(P_{4})}$v4t(Ours) 41.0 73.6 25.5 79.1 50.3 25.7 27.8 21.0 28.4 41.4 46.5
${}^{(P_{5})}$v5s(Ours) 47.4 64.3 27.2 56.5 69.9 23.1 31.6 25.7 29.9 38.3 37.7
${}^{(P_{6})}$v7t(Ours) 35.0 64.3 23.0 56.7 44.5 33.5 31.6 25.3 26.1 38.9 47.7
${}^{(P_{7})}$v10s(Ours) 35.3 62.0 23.4 59.9 45.7 24.2 66.1 27.2 28.9 40.3 45.0
${}^{(P_{8})}$v12s(Ours) 44.0 57.3 37.9 55.1 45.1 22.2 47.7 63.7 32.4 27.4 42.7
${}^{(P_{9})}$(Ours) 42.7 62.4 29.2 60.8 50.7 24.3 28.8 24.8 38.5 43.7 45.5
${}^{(P_{10})}$(Ours) 64.2 79.5 43.1 75.2 62.7 27.1 41.2 31.8 37.0 46.5 52.0
${}^{(P_{11})}$-DETR(Ours) 36.9 55.2 22.2 48.5 40.1 22.3 26.7 23.5 27.1 37.9 56.3
${}^{(P_{j})}$ource 29.7 38.1 18.5 36.9 35.0 18.4 16.0 17.4 23.3 33.6 27.3
       Gray 16.5 25.2 12.8 30.0 15.5 13.5 13.2 12.3 17.3 34.4 16.5
${}^{(P_{m})}$andom 16.9 22.2 12.2 26.9 15.5 12.3 11.2 12.7 17.6 31.7 15.6
${}^{(P_{n})}$andom 17.0 19.0 10.4 20.7 11.7 12.0 9.8 10.6 16.4 25.9 15.5
$P_{1}$      $P_{2}$      $P_{3}$      $P_{4}$      $P_{5}$      $P_{6}$      $P_{A}$      $P_{B}$      $P_{C}$      $P_{m}$
[Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]      [Uncaptioned image]
$P_{7}$      $P_{8}$      $P_{9}$      $P_{10}$      $P_{11}$      $P_{j}$      $P_{D}$      $P_{E}$      $P_{F}$      $P_{n}$
@@ 4.1 Experimental Setup
In our experiments, we train and evaluate the adversarial patches on the INRIA person dataset [12], which consists of 614 training images and 288 test images. To meet the input requirements of the detector, we pad all images with gray pixels to make them square and then resize them to 640 $\times$ 640. Furthermore, we evaluate the transferability of different adversarial patch methods on the MPII Human Pose dataset [2], COCO dataset [35] and 1,000 images from diverse scenarios collected from the Internet.
Victim Models
Our target a variety of detectors, including YOLOv3, YOLOv3-tiny [43], YOLOv4, YOLOv4-tiny [5], YOLOv5s [28], YOLOv7-tiny [53], YOLOv10s [52], YOLOv12s [51], FasterRCNN [44] and DETR [7] and RT-DETR [60]. We will abbreviate YOLO as YL and FasterRCNN as FRCNN for ease of presentation, such as YOLOv4-tiny as YLv4t. All models were pre-trained on the COCO dataset [35].
@@
We treat the person as the target object class, setting the batch size to 32, the maximum number of epochs $N_{iter}=200$, DDIM sampling steps $T=50$, unconditional embedding optimization iterations $N_{u}=10$, learning rates $\eta_{u}=0.01$ and $\eta_{p}=0.003$, and perturbation constraint $\epsilon=0.5$. We use the Adam optimizer [29] with its default parameters. Textual prompts $\mathcal{P}$ are automatically generated using BLIP [33]. The adopted version of Stable Diffusion [45] is v1.4. Since we generate irregular patches, we set the patch size to match the pixel area of a square patch with a scale of $\tau=0.2$. To achieve stronger performance, we apply two iterations of optimization to the patch using DiffPatch. We evaluate the effectiveness of different patch generation methods using both the Attack Success Rate (ASR) and Average Precision (AP). The ASR measures the success of the patches in misleading the detector, while AP reflects the overall decline in detection accuracy. The confidence threshold and IoU threshold for the detector are both set to 0.5. All experiments were conducted on a single NVIDIA A100 GPU.
@@ 4.2 Attack Performance Evaluation
We generate a series of adversarial patches targeting different models using DiffPatch and compare them on the INRIA dataset with existing methods, including AdvYL [50], T-SEA [25], DAP [20], AdvART [19], NPAP [24], and D2D [34]. The detailed results are reported in Table 1. As shown, DiffPatch ($P_{1}-P_{11}$) achieves the highest Attack Success Rate (ASR), improving by 10.1% over AdvYL on YOLOv4-tiny. When compared to naturalistic adversarial patches ($P_{C}$ and $P_{D}$), our patch trained on YOLOv4 results in a 0.2% and 4.1% ASR increase, respectively. Additionally, as demonstrated in the Appendix B, DiffPatch exhibits strong performance in the AP metric, performing comparably to unnatural patch generation methods.
@@ 4.3 Cross-dataset Evaluation
Given the relatively simple scenes in the INRIA dataset, we further evaluate the attack performance of the patches in more diverse environments. We train patches using YOLOv4 with various methods and test them across multiple datasets, including MPII and COCO. Due to the large size of these datasets, we randomly sample a subset of images containing people for testing, specifically 729 images from COCO and 569 images from MPII. Additionally, since the images from these two datasets are relatively outdated and of low resolution, we also collect 1,000 high-resolution images from the Internet, representing diverse scenes (denoted as ”Collected”). As shown in Table 2, our method significantly reduces the AP while substantially increasing the ASR. On average, the ASR shows a 17.5% improvement over the unnatural patch method T-SEA and a 2.6% improvement over the natural patch method NPAP. Several example images and their detection results with adversarial patches are provided in the Appendix A.
@@
Table 2: Attack performance of different methods across datasets.
Ensemble Attack
We also evaluate an ensemble version of DiffPatch, which trains the patch using averaged gradients from multiple models. We compare this with two other methods, NPAP and D2D, where NPAP was trained on YOLOv3 and YOLOv4, and D2D was trained on all considered models. Our ensemble DiffPatch was similarly trained on all models. The results, presented in the Appendix C, demonstrate that our method outperforms NPAP and D2D by a significant margin.
@@
$Patch$
20406080Average ASR(%)406080Average CLIP Sim.(%)ASRCLIP Sim.
Figure 4: The effect of optimization constraint $\epsilon$ on latent vectors.
@@ 4.4 Ablation Studies
Optimization Constraint
To ensure the natural appearance of the generated patches, we constrain each optimization step using the $L_{\infty}$ norm. The choice of $\epsilon$ will influence visual naturalness of the generated patch and its attack performance. To quantify the natural appearance of the patches, we use GPT-4o [1] to generate descriptive text for reference images (denoted by ’-’) and compute the CLIP similarity [41] between these descriptions and the generated patches. The corresponding descriptions can be found in the Appendix D. As illustrated in Figure 4, we obtain patches with varying styles under different constraints using DiffPatch and evaluate both the ASR and CLIP similarity for each. Based on the results, we select $\epsilon=0.5$ as the optimal value, as it achieves a favorable balance between attack performance and natural appearance.
@@
Here, we compare our IoU-Detection Loss with the Common Detection Loss under the same patch training setting. As shown in Figure 5, compared to the Common Detection Loss, our IoU-Detection Loss demonstrates greater stability during training process and achieves higher attack performance in an earlier time.
@@
Here, we investigate the impact of patch shapes, as used for $P_{m}$ and $P_{n}$ in Table 1. We compare the performance of random noise images with square and irregular shapes (of the same size as DiffPatch). From the ASR and AP results in Table 1, it is evident that square patches yield higher attack effectiveness compared to irregular patches. In other words, designing irregular patches with high attack success rates is more challenging.
@@
Here, we evaluate the results of latent vectors at different timestamps, assessing both the attack performance and the similarity to textual descriptions for patches generated from diffusion trajectories of varying lengths. As shown in Figure 6, starting diffusion optimization from $\frac{T}{2}$ (25 steps) not only achieves a high success rate but also produces images that align more closely with textual descriptions, thereby maintaining greater semantic consistency with the original images.
@@
The target mask mechanism helps mitigate semantic loss in images to some extent. Here, we compare the training results with and without the use of this mechanism. Without the target mask, the patches exhibit quality issues during optimization, as illustrated in Figure 7.
@@
As shown in Table 3, iterative optimization can further enhance the adversarial strength of the patch. However, it also tends to reduce the natural appearance of the patch.
@@
4.5 Adversarial T-shirt Dataset
To better evaluate the effectiveness of our attack and create a valuable physical-world adversarial patch dataset for future research, we designed 9 physical-world adversarial T-shirts using our generated patches. We printed these patches on T-shirts and captured photos in a variety of indoor and outdoor scenarios, such as laboratory, campus, cafeteria, subway station, and shopping mall, with the assistance of 20 participants (with ethics approval).
In total, we collected 1,131 images, including both individual and group photos (ranging from 2 to 10+ persons). We name the dataset AdvT-shirt-1K, which includes detailed annotations for person and patch locations (bounding boxes). The collection and annotation process of the dataset spanned six months, which involves the refinement of our DiffPatch method for multiple times. The statistics of the dataset is shown in Figure 8 with several example images are provided in Figure 1. The results shown in the right subfigure of Figure 8 demonstrate the effectiveness of the adversarial T-shirts in evading the YOLOv5s detector. Although small in scale, our dataset provides a practical resource for developing and testing the real-world effectiveness of adversarial patch detectors.
Figure 7: The impact of target mask on adversarial patch optimization. Without target mask control, adversarial patches exhibit noticeable detail distortions (indicated by the red boxes).
Figure 8: Left: The distribution of the images in AdvT-shirt-1K. Right: The detection performance of YOLOv5s on AdvT-shirt-1K under different confidence thresholds.